In buying a computer disk, the buyer was offered the options of paying 250 cash at the end of 30 days or 270 at the end of 120 days. At what rate is the buyer paying simple interest if he agrees to pay at the end of 120 days

1
Expert's answer
2022-01-20T11:45:51-0500

In buying a computer disk, the buyer was offered the options of paying 250 cash at the end of 30 days or 270 at the end of 120 days. At what rate is the buyer paying simple interest if he agrees to pay at the end of 120 days


Answer

"SI=\frac{P×R×T}{100}"

"SI=A-P"

"A=270"

"P=250"

"SI=270-250=20"

"T=120-30=90" months

"T=\frac{1}{4}" an year


"R=\frac{SI×100}{P×T}"

"R=\frac{20×100×4}{250×1}"

"=32%"%